An old human skull was unearthed excavated at ogbomoso and its 14-6C/12-6C ratio was only 0.326 of that in the atmosphere today. How old is the skull, if the half-life of 14-6C is 5700 years
1
Expert's answer
2020-10-12T07:45:26-0400
"N=N_0\cdot 2^{-\frac{t}{T}}\\0.326= 2^{-\frac{t}{5700}}\\t=9217\ yr"